OpenStudy (anonymous):

Carbon-14 has a half life of 5,730 years. If a sample contains 70 mg originally, how much is left after 17,190 years? @beccaboo333

OpenStudy (beccaboo333):

What do you think?

OpenStudy (anonymous):

I dont know. I do not understand.

OpenStudy (beccaboo333):

Give me a moment

OpenStudy (beccaboo333):

There is 3 half-life time in 17190 years. So this, ( 17190/5730 = 3) If to use the half life formula you'd have, y=original amount * (1/2)^x y=70(1/2)^3 y=8.75 mg Welcome to OpenStudy.